OpenStudy (anonymous):

solve the system y=2-x and y=4x+7

OpenStudy (jdoe0001):

\(\bf {\color{red}{ y}}=2-x\\ y=4x+7 \\ \quad \\ y=4x+7\implies y=4({\color{red}{ 2-x}})+7\) solve for "y" then substitute that value in either equation and solve for "x"

OpenStudy (jdoe0001):

hmmm lemme redo that =)

OpenStudy (jdoe0001):

\(\bf {\color{red}{ y}}=2-x\\ y=4x+7 \\ \quad \\ y=4x+7\implies ({\color{red}{ 2-x}})=4x+7\) solve for "x" rather, and then substitute that on either equation and solve for "y"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it will be (-1,3) or (1,3)

OpenStudy (jdoe0001):

what did you get for "x"?

OpenStudy (anonymous):

4x^2 I thank I gat confused with math

OpenStudy (jdoe0001):

hmmm tis just a linear equation -> \(\bf ({\color{red}{ 2-x}})=4x+7\implies 2-x=4x+7\)

OpenStudy (anonymous):

thank you that helped a lot

OpenStudy (jdoe0001):

yw